OpenStudy (anonymous):

Which expression is a fourth root of (-1 + iSqrt3) A.4sqrt2(cos(280)+ isin(280)) B. 4sqrt2(cos(30)+ isin(30)) C.4sqrt2(cos(60)+ isin(60)) D.4sqrt2(cos(320)+ isin(320))

OpenStudy (anonymous):

\[(-1+ i \sqrt3) \]

hartnn (hartnn):

find \(\theta = \tan^{-1} (-\sqrt 3) = ...\)

OpenStudy (anonymous):

-60

OpenStudy (anonymous):

so should i be getting \[\sqrt[4]{2}(\cos(60) + isin(60)\]

OpenStudy (anonymous):

that was wrong but thats okay XD

hartnn (hartnn):

sorry for the late reply. since (-1 + iSqrt3) is in 2nd quadrant, theta will be 120 degrees so we have r<theta = 2 <120 taking 4th root of complex number means taking 4th root of 'r' and dividing the angle theta by 4 so, 4th root will be \(\sqrt[4]2 \angle 120/4 = \sqrt [4]2 \angle 30 =\sqrt [4]2(\cos 30 +i \sin 30) \)

OpenStudy (anonymous):

no problem :) thank you again, your help is very Super Saiyan worthy!

hartnn (hartnn):

welcome ^_^ i'll take the compliment, thanks ;)